OpenStudy (anonymous):

If the population of Town B is 50% greater than the population of Town A and the populations of Town C is 20% greater than the population of Town A, then what percent greater is the population of Town B tan the population of Town C?

OpenStudy (anonymous):

Can you explain?

OpenStudy (blacksteel):

Let A = the population of town A, and B and C be defined similarly. Then B = 1.5A C = 1.2A To find the percentage size of B compared to C, divide the population of B by C. B/C = 1.5A/1.2A = 1.5/1.2 = 1.25, so B is 125% of C, or 25% greater.

Directrix (directrix):

@sakigirl This looks as if it may be an SAT question. If so, you can make up friendly numbers to get the answer. Town A: 100 people Town B: 150 people which is 50% greater population than Town A Town C: 120 people which is 20% greater population than Town A -------- Town B: 150 people Town C: 120 people 150/120 = 15/12 = 5/4 = 1.25 which is 25% larger as computed by @Blacksteel